A 16-liter radiator is filled with a solution of 40% antifreeze. How much should you drain from the radiator and replace with pure antifreeze to obtain a 60% antifreeze solution?
Answer by stanbon(75887)

A 16-liter radiator is filled with a solution of 40% antifreeze. How much should you drain from the radiator and replace with pure antifreeze to obtain a 60% antifreeze solution?
--------------------------------
Antifreeze = antifreeze
0.4*16 - 0.4x + x = 0.6*16
--------
Multiply thru by 10 to get:
4*16 - 4x + 10x = 6*16
6x = 2*16
x = 32/6
x = 5 1/3 liters is the amount to be drained out and replaced
=======================
Cheers,
Stan H.